# T-GVC: Trajectory-Guided Generative Video Coding at Ultra-Low Bitrates

### v3

@@ 1 Introduction
One of the core challenges in video coding lies in effectively modeling inter-frame dependencies to reduce temporal redundancy and enhance coding efficiency. While traditional video coding standards (e.g., H.266/VVC (Bross et al. 2021) and AVS3 (Zhang et al. 2019)) have achieved remarkable progress, their reliance on handcrafted motion compensation and transform-quantization modules exhibits limitations in modeling non-rigid motions and nonlocal spatio-temporal dependencies, particularly under Ultra-Low Bitrate (ULB) scenarios with constrained bandwidth. Benefiting from end-to-end optimization strategy, recently emerging deep learning-based video coding approaches (Man et al. 2024; Li, Li, and Lu 2023, 2024; Jiang et al. 2024) have demonstrated promising potential to surpass conventional schemes. However, these methods primarily emphasize pixel-level signal fidelity and suffer from semantic information loss in ULB scenarios due to limited temporal context information. Although recent attempts (Yang, Timofte, and Van Gool 2022; Du, Liu, and Ling 2024; Du et al. 2022) have leveraged GAN (Goodfellow et al. 2020) to improve the perceptual quality, they still adopt similar frameworks as (Li, Li, and Lu 2023, 2024; Jiang et al. 2024; Yang et al. 2020) to model temporal contextual information, which constraints their ability to achieve lower bitrates. Consequently, preserving critical semantic information while enhancing perceptual quality has become a pressing challenge for ULB video compression.
@@
Recent breakthroughs in video generation technology have provided new feasibility for video coding under ULB conditions, catalyzing the emergence of the concept of ”generative video coding” (Chen et al. 2024b). Unlike generation tasks in computer vision, generative video coding primarily aims to achieve content-faithful reconstruction under strict bitrate constraints by leveraging the strong priors of generative models as well as compact spatio-temporal guidance. Despite these theoretical advantages, most existing generative video coding schemes are tailored to specific types of videos, such as human face video (Chen et al. 2024a), human body video (Wang et al. 2022, 2023), and small motion video (Yin et al. 2024), due to the limited capabilities of earlier-stage generative models.
@@
The rapid advancement of video diffusion models (VDMs) (Xing et al. 2024b) has significantly intensified research interest in extending generative video coding to natural video content. Specifically, some recent works (Zhang et al. 2024b; Wan, Zheng, and Fan 2025) attempt to model spatio-temporal information solely based on texts and keyframes from the original video. However, although text serves as an effective modality for representing high-level semantics of video, text-guided generative models often encounter difficulties in reconstructing motion with high perceptual quality and may even produce visibly inferior results in certain cases. In parallel, image-to-video or image-text-to-video models (Xing et al. 2024a; Zhang et al. 2024a) have shown the potential to synthesize motion sequences conditioned on the structural and textual information derived from a pair of keyframes. Nevertheless, these methods still struggle to ensure temporal consistency and visual stability when applied to complex real-world scenarios. The generated outputs may exhibit unexpected artifacts, such as unnatural motion patterns or missing dynamic objects, compromising the temporal semantic fidelity with the original video. To allow for more precise generation, an intuitive idea is to take advantage of low-level motion information, as shown in Figure 1. Unfortunately, employing dense motion representation incurs a higher bitrate cost for VDMs to generate high-fidelity videos. It is crucial to develop an effective strategy to extract compact motion guidance that preserves key temporal semantic information from the original video.
@@
Another important aspect concerns the conditional guidance mechanisms employed in the generative model. Most existing controllable video generation methods utilize ’classifier-free guidance’ (Ho and Salimans 2022) to fine-tune pre-trained video diffusion models, such as Stable Video Diffusion (SVD) (Blattmann et al. 2023) and VideoCraft (Chen et al. 2023b, 2024c), or adopt a ControlNet-like adapter (Zhang, Rao, and Agrawala 2023; Niu et al. 2024) to steer the generation process according to user specifications. However, the loss of critical structural and textural details in compressed keyframes significantly impairs the model’s ability to reconstruct motion-aligned videos, particularly under ULB conditions. Moreover, these pre-trained generative models suffer from limited adaptability in the context of video coding tasks, as the compact temporal guidance (e.g., motion trajectory) may deviate substantially from the original conditioning domains. How to effectively leverage compact guidance information to control the generative model in reconstructing high-fidelity and motion-aligned videos also requires further explored.
In response to the critical challenge of generative video coding under ULB conditions, we propose a Trajectory-guided Generative Video Coding framework. To balance the trade-off between temporal semantic information preservation and bitrate saving, a semantic-aware sparse motion sampling pipeline is first proposed to bridge low-level motion tracking with high-level semantic understanding. Specifically, we track the motion of pixels on a pre-defined grid and classify them into distinct motion instances according to motion patterns. To further improve coding efficiency, a sparse sampling of motion instances is performed based on their semantic importance, representing motion as a set of trajectory points. The variations in trajectory points reflect different motion patterns (such as translation, occlusion, deformation, etc.), which preserves critical temporal semantic information of the original video while significantly reducing the bitrate. To guide the diffusion-based generative model in reconstructing motion-aligned videos without compromising its inherent generative capabilities, we propose a training-free guidance approach. In contrast to existing methods that constrain on intermediate feature maps, our approach directly imposes constraints on the latent space of diffusion model via a lightweight yet effective guidance function, ensuring that the overall trajectories of target motion instances align with real-world motion paths.
@@
  • We propose a semantic-aware sparse motion sampling pipe-line tailored for generative video coding, which bridges low-level motion tracking with high-level semantic understanding by encoding motion as sparse trajectory points, significantly reducing bitrate while preserving temporal semantics.
@@ 2.1 Generative video coding
Most end-to-end video coding frameworks optimized for signal fidelity (Shi and Lu 2024) face critical challenges in maintaining satisfactory perceptual quality under ULB constraints. Recently, generative approaches have been explored to improve perceptual quality in ULB scenarios, giving rise to the concept of generative video coding (Chen et al. 2024b). Early generative video coding methods primarily targeted specific content types, such as facial or human-centric videos. For instanse, Wang et al. (Wang, Mallya, and Liu 2021) first proposed video synthesis based on reference frames and keypoints for facial video compression. Subsequently, Feng et al. (Feng et al. 2021) incorporated adversarial training with additional modalities to improve visual fidelity, and such a strategy was also adopted in (Oquab et al. 2021; Hong et al. 2022). In (Shukor et al. 2022), Shukor et al. projected video frames into the latent space of StyleGAN (Richardson et al. 2021) for efficient compression, while Chen et al. (Chen et al. 2023a) introduced a dynamic reference refresh mechanism to enhance temporal stability. Beyond facial contents, Wang et al. (Wang et al. 2022) decomposed human-centric videos into texture and structure via contrastive learning and human pose keypoints, and future work (Wang et al. 2023) leveraged Principal Component Analysis (PCA) to compress motion features for more compact human video representations.
For natural videos, early approaches (Yang, Timofte, and Van Gool 2022; Du et al. 2022; Du, Liu, and Ling 2024) employed GANs to enhance perceptual quality under low bitrate conditions. More recently, a framework based on VQ-VAE (Qi et al. 2025) further performed transform coding in the latent space to achieve high realism and high fidelity at ULB. However, constrained by the inherent architectures of conventional frameworks (Li, Li, and Lu 2023, 2024; Jiang et al. 2024; Yang et al. 2020), these methods are difficult to generalize to even lower bitrate scenarios. Inspired by Cross-Modal Compression (CMC) in the field of image compression (Li et al. 2021; Zhang et al. 2023; Gao et al. 2023, 2024), Zhang et al. (Zhang et al. 2024b) designed Cross-Modal Video Compression (CMVC) to explore the potential of applying multimodal large language models for ULB video compression. In CMVC, a video is segmented into multiple semantically coherent clips, each of which is decomposed into content (a keyframe) and motion (text descriptions). On the decoder side, keyframes are reconstructed first, followed by the generation of intermediate frames guided by the extracted ’motion’ information. Similarly, Wan et al. (Wan, Zheng, and Fan 2025) utilized a text-guided video diffusion model for video reconstruction, incorporating a more elaborate keyframe selection strategy. While these approaches are able to preserve essential semantic contents at ULB, they still suffer from issues with temporal consistency when handling complex motion patterns. Overall, despite significant progress in generative video coding for specific content types, extending these techniques to natural video, particularly in dynamic and motion-rich scenes, remains an open and challenging topic.
@@ 2.2 Motion-conditioned diffusion model
Text-conditioned video diffusion models (VDMs) have demonstrated remarkable capabilities in synthesizing visually striking content. However, they still struggle to achieve precise outcome control, frequently generating unrealistic sequences that violate physical plausibility. To address this limitation, recent studies have increasingly focused on incorporating explicit motion guidance into diffusion models to improve temporal coherence and realism of generated videos. As one of the pioneering works, MCDiff (Chen et al. 2023c) performed video synthesis conditioned on motion signals by providing the first video frame along with a sequence of stroke motions. A flow completion model was employed to convert sparse inputs into dense motion maps, followed by an auto-regressive strategy to predict subsequent frames. Building upon this idea, DragNUWA (Yin et al. 2023) improved the motion-guided scheme by integrating text, image, and trajectory information into VDMs, enabling more fine-grained control over video content generation. More recently, MotionCtrl (Wang et al. 2024) proposed multiple dedicated modules to manage camera motion and object motion independently, while Tora (Zhang et al. 2024c) first integrated trajectories into Diffusion Transformer (DiT) framework (Peebles and Xie 2023), producing longer videos with better physical realism.
All of the aforementioned guidance approaches can be seamlessly compatible with existing diffusion-based video generation frameworks, enhancing the overall motion modeling capabilities. However, they typically require training new diffusion models or motion control modules from scratch. To mitigate the trade-off between quality and computation cost, more recent studies (Xiao et al. 2024; Zhang et al. 2025) have explored the use of classifier guidance (Dhariwal and Nichol 2021) to control the video generation process in a more efficient manner. Leveraging a PCA-based analysis of VDMs, Xiao et al. (Xiao et al. 2024) revealed the presence of intrinsic motion-aware features within the latent representations of VDMs, and proposed a training-free framework for motion control. Similarly, Zhang et al. (Zhang et al. 2025) extracted motion patterns as soft trajectories from the feature maps of the temporal attention modules in VDMs, and performed classifier guidance (Dhariwal and Nichol 2021) to align the denoising process with the target motion patterns. These approaches have shown the flexibility and effectiveness of the training-free video motion control framework, which provided valuable inspiration for our work.
@@ 3 METHODOLOGY
3.1 Overview
The overall structure of the proposed T-GVC framework is shown in Figure 2, which follows the common paradigm of generative video coding. Within the framework, the semantic-aware sparse motion sampling pipeline serves as a sparse motion sampler to capture motion from inter-frames, while the latent space guidance mechanism plays a role in motion correction during the DDIM (Song, Meng, and Ermon 2020) denoising process. In the remainder of this section, we will describe the entire coding pipeline of T-GVC and elaborate on each module. For ease of understanding, we list the main symbols used in this paper in Table 1.
@@ 3.2 Keyframe selection and compression
Following (Zhang et al. 2024b), we first segment the source video into multiple clips and select the frame with the smallest pairwise semantic similarity as the keyframe, which is measured via a CLIP-based criteria (Radford et al. 2021). To prevent visual artifacts caused by abrupt scene transitions, we additionally introduce a semantic-aware scene cut detection mechanism based on inter-frame similarity analysis. Specifically, when the semantic similarity score between two consecutive frames falls below a pre-defined threshold, both frames are designated as keyframes to ensure seamless transitions. Therefore, more than one keyframes may be selected within each video clip.
Each pair of consecutive keyframes $\{K_{i},K_{i+1}\}$, together with their associated intermediate frames, constitutes a new clip $\{F_{0},F_{1},...,F_{L-1}\}$. The clip structure serves as the fundamental processing unit for subsequent motion sampling and semantic-consistent video reconstruction. Notably, since keyframes carry critical semantic information, their reconstruction on the decoder side must maintain high perceptual quality even under stringent bitrate constraints, thereby preserving essential textural details and structural features to facilitate subsequent intermediate frame reconstruction. As this work primarily focuses on inter coding, we adopt the off-the-shelf MS-ILLM (Muckley et al. 2023) for keyframe compression, owing to its demonstrated ability to preserve superior perceptual fidelity under aggressive bitrate constraints.
@@ 3.3 Semantic-aware sparse motion sampling
Given two keyframes $\{K_{i},K_{i+1}\}$ and their intermediate frames, we design a bidirectional tracking scheme to extract dense motion trajectories with improved temporal consistency. For each clip $\{F_{0},F_{1},...,F_{L-1}\}$ with temporal length $L$, we initialize isometric grids (grid size $N_{grid}\times N_{grid}$) in the first frame and employ Co-Tracker (Karaev et al. 2024) to trace pixel-wise displacements originating from the grid vertices. The generated forward trajectory sequences over $L$ frames are represented as temporal chains of 3D coordinate points:
@@
To mitigate trajectory loss caused by occlusions or rotation, we reverse the temporal order of each clip and extract the corresponding backward trajectory sequences using the same procedure:
$\mathcal{T}_{bwd}=\left\{\left.T_{j}^{bwd}\right|T_{j}^{bwd}={(x_{t}^{j},y_{t}^{j},V_{t}^{j})}_{t=0}^{L-1}\right\}$ (2)
Subsequently, the backward trajectories are temporally realigned and fused with their forward counterparts to synthesize the final dense trajectories, which holistically preserve the motion characteristics between keyframes:
@@
To differentiate between distinct motion patterns, the raw trajectories are clustered using HDBSCAN (McInnes et al. 2017), a hierarchical density-based algorithm particularly effective in handling variable cluster densities and suppressing noise, which is critical for processing real-world videos with irregular motions. In this stage, $T_{i}$ is first transformed into 2D trajectories by excluding points where $V_{t}^{i\ }=0$. Then, we formulate each $T_{i}$ as a spatio-temporal feature vector $f^{i}_{traj}$ as follows:
[…]
By clustering these features, the corresponding trajectories are categorized into different motion instances $\mathcal{T}^{0},\mathcal{T}^{1},...,\mathcal{T}^{n-1}$ and form the corresponding motion masks, as illustrated in Figure 2.
@@
The extracted dense motion trajectories inherently encapsulate the spatio-temporal dynamics of the source video, enabling the decoder to reconstruct inter-frame semantic relationships through diffusion-based generation (Section 3.4). However, naively encoding these raw trajectories introduces significant bitrate overhead. To reduce inter-frame bitrate while preserving semantically critical motion information, we propose a trajectory sampling strategy driven by semantic loss awareness.
@@
We select the motion whose inter-frame semantic importance exceeds the threshold and extract sparse trajectories $\tau^{i}$ through k-means clustering, where the keypoint quantity $K$ for each instance is determined by the semantic importance score $S^{i}_{inter}$ and the quantity of trajectories $N^{i}$:

### v4

@@ 1 Introduction
One of the core challenges in video coding lies in effectively modeling inter-frame dependencies to reduce temporal redundancy and enhance coding efficiency. While traditional video coding standards (e.g., H.266/VVC (bross2021overview) and AVS3 (zhang2019recent)) have achieved remarkable progress, their reliance on handcrafted motion compensation and transform-quantization modules exhibits limitations in modeling non-rigid motions and nonlocal spatio-temporal dependencies, particularly under Ultra-Low Bitrate (ULB) scenarios with constrained bandwidth. Benefiting from the development of deep learning, recently emerging neural network-based video coding approaches (10155660; li2024neural; jiang2024ecvc; Man2024; li2024object; Man2025; tang2025neural) have demonstrated promising potential to surpass conventional schemes. However, these methods primarily emphasize pixel-level signal fidelity and suffer from semantic information loss in ULB scenarios due to limited temporal context information. Although recent attempts (yang2022perceptual; du2024cgvc; du2022generative) have leveraged GAN (goodfellow2020generative) to improve the perceptual quality, they still adopt similar frameworks as (li2023neural; li2024neural; yang2020rlvc) to model temporal contextual information, which constraints their ability to achieve lower bitrates. Consequently, preserving critical semantic information while enhancing perceptual quality has become a pressing challenge for ULB video compression.
@@
Recent breakthroughs in video generation technology have provided new feasibility for video coding under ULB conditions, catalyzing the emergence of the concept of “generative video coding” (chen2024generative). Unlike generation tasks in computer vision, generative video coding primarily aims to achieve content-faithful reconstruction under strict bitrate constraints by leveraging the strong priors of generative models as well as compact spatio-temporal guidance. Despite these theoretical advantages, most existing generative video coding schemes are tailored to specific types of videos, such as human face video (chen2024generative2), human body video (wang2023extreme; chen2025rethinking), and small motion video (yin2024compressing), due to the limited capabilities of earlier-stage generative models.
@@
The rapid advancement of video diffusion models (VDMs) (xing2024survey) has significantly intensified research interest in extending generative video coding to natural video content. Specifically, some recent works (zhang2024video; wan2025m3) attempt to model spatio-temporal information based on texts and keyframes, showing the potential to synthesize motion sequences conditioned on sparse and high-level information derived from the original video. Nevertheless, these methods still struggle to ensure temporal consistency and visual stability when applied to complex real-world scenarios. The generated outputs may exhibit unexpected artifacts, such as unnatural motion patterns or missing dynamic objects, compromising the temporal semantic fidelity with respect to the original video. To enable more precise natural video generation, an intuitive idea is to take advantage of low-level motion information, as illustrated in Figure 1. Unfortunately, employing dense motion representation incurs a higher bitrate overhead for VDMs to synthesize high-fidelity videos. It is crucial to develop an effective strategy to extract compact motion guidance that preserves key temporal semantic information from the original video.
Another important aspect concerns the conditional guidance mechanisms employed in the generative model. Most existing controllable video generation methods utilize ‘classifier-free guidance’ (ho2022classifier) to fine-tune pre-trained video diffusion models, such as Stable Video Diffusion (SVD) (blattmann2023stable) and VideoCraft (chen2023videocrafter1; chen2024videocrafter2), or adopt a ControlNet-like adapter (zhang2023adding; niu2024mofa; liu2025sketchvideo) to steer the generation process according to user specifications. However, the loss of critical structural and textural details in compressed keyframes severely impairs the model’s ability to reconstruct motion-aligned videos, particularly under ULB conditions. Moreover, these pre-trained generative models suffer from limited adaptability in the context of video coding tasks, as the compact temporal guidance (e.g., motion trajectory) may deviate substantially from the original conditioning domains. How to effectively leverage compact guidance information to control the generative model in reconstructing high-fidelity and motion-aligned videos also requires further exploration.
@@
In response to the critical challenge of generative video coding under ULB conditions, we propose a Trajectory-Guided Generative Video Coding framework. To balance the trade-off between temporal semantic information preservation and bitrate saving, a semantic-aware sparse motion sampling pipeline is first proposed to bridge low-level motion tracking with high-level semantic understanding. Specifically, we track the pixel-wise motion on a pre-defined grid and classify them into distinct motion instances according to motion patterns. To further improve coding efficiency, a sparse sampling of motion instances is performed based on their semantic importance, representing motion as a set of sparse trajectory points. The variations in trajectory points capture diverse motion patterns (such as translation, occlusion, deformation, etc.), which preserves critical temporal semantic information of the original video while significantly reducing the bitrate. To guide the diffusion-based generative model in reconstructing motion-aligned videos without compromising its inherent generative capabilities, we propose a training-free guidance approach. In contrast to existing methods that constrain on intermediate feature maps, our approach directly imposes constraints on the latent space of diffusion model via a lightweight yet effective guidance function, ensuring that the overall trajectories of target motion instances align with real-world motion paths.
@@
  • We propose a semantic-aware sparse motion sampling pipeline tailored for generative video coding, which bridges low-level motion tracking with high-level semantic understanding by encoding motion as sparse trajectory points, significantly reducing bitrate while preserving temporal semantics.
@@ 2.1 Generative video coding
To ensure satisfactory perceptual quality, generative video coding has been explored in Ultra-Low Bitrate (ULB) scenarios. Early methods in this domain mainly focused on specific content types, such as facial videos (chen2024generative2) or human-centric videos (wang2022disentangled; wang2023extreme). For natural videos, some approaches (yang2022perceptual; du2022generative; du2024cgvc) employed GANs to enhance perceptual quality under low bitrate conditions. A more recent framework based on VQ-VAE (qi2025generative) further performed transform coding in the latent space to achieve high realism and high fidelity at ULB. However, constrained by the inherent architectures of conventional frameworks (yang2020rlvc; li2023neural; li2024neural), these methods are difficult to generalize to even lower bitrate scenarios.
Inspired by Cross-Modal Compression (CMC) in the field of image compression (li2021cross; zhang2023rethinking; gao2023cross; gao2024rate), zhang2024video designed Cross-Modal Video Compression (CMVC) to explore the potential of applying multimodal large language models for ULB video compression. In CMVC, a video is segmented into multiple semantically coherent clips, each of which is decomposed into content (a keyframe) and motion (text descriptions)(10003241). On the decoder side, keyframes are reconstructed first, followed by the generation of intermediate frames guided by the extracted ‘motion’ information. Similarly, wan2025m3 utilized a text-guided video diffusion model for video reconstruction, incorporating a more elaborate keyframe selection strategy. While these approaches are able to preserve essential semantic contents at ULB, they still suffer from issues with temporal consistency when handling complex motion patterns. Overall, despite significant progress in generative video coding for specific content types, extending these techniques to natural video, particularly in dynamic and motion-rich scenes, remains an open and challenging topic.
@@ 2.2 Motion-conditioned diffusion model
Text-conditioned video diffusion models (VDMs) have demonstrated remarkable capabilities in synthesizing visually striking content. However, they still struggle to achieve precise outcome control, frequently generating unrealistic sequences that violate physical plausibility. To improve temporal coherence and realism, recent efforts introduced motion guidance into VDMs. Early works like MCDiff (chen2023motion) and DragNUWA (yin2023dragnuwa) incorporated sparse motions and multimodal inputs, while later models such as MotionCtrl (wang2024motionctrl) and Tora (zhang2024tora; zhang2025tora2) introduced finer-grained control via explicit trajectory modeling. However, they typically require training new diffusion models or motion control modules from scratch. To mitigate the trade-off between quality and computation cost, more recent studies (xiao2024video; zhang2025training) have explored the use of classifier guidance (dhariwal2021diffusion) to control the video generation process in a more efficient manner, revealing the presence of intrinsic motion-aware features within the latent representations of VDMs. These approaches have shown the flexibility and effectiveness of the training-free video motion control framework, which provided valuable inspiration for our work.
@@ 3 Methodology
The overall structure of the proposed T-GVC framework is shown in Figure 2, which follows the common paradigm of generative video coding. In the remainder of this section, we will describe the entire coding pipeline of T-GVC and elaborate on each module with key symbols summarized in Table 1.
@@ 3.1 Keyframe selection and compression
Following (zhang2024video; 11146420), we segment the source video into clips and select keyframes based on minimal pairwise semantic similarity, measured via a CLIP-based criteria (radford2021learning). To robustly handle scene transitions, we further incorporate a semantic-aware scene cut detection mechanism based on inter-frame similarity analysis. Specifically, when the semantic similarity between two consecutive frames falls below a pre-defined threshold, both frames are designated as keyframes to ensure seamless scene transitions. Therefore, more than one keyframes may be selected within each video clip. Each pair of consecutive keyframes $\{K_{i},K_{i+1}\}$, along with the associated intermediate frames, constitutes a new clip $\{F_{0},F_{1},...,F_{L-1}\}$. The clip structure serves as the fundamental processing unit for subsequent motion sampling and semantic-consistent video reconstruction.
Notably, since keyframes encapsulate critical semantic information, their reconstruction on the decoder side must maintain high perceptual quality even under stringent bitrate constraints, thereby preserving essential textural details and structural features to facilitate subsequent intermediate frame reconstruction. As this work primarily focuses on inter-frame coding, we adopt the off-the-shelf MS-ILLM (muckley2023improving) for keyframe compression, owing to its demonstrated ability to preserve superior perceptual fidelity under aggressive bitrate constraints.
@@ 3.2 Semantic-aware sparse motion sampling
Given two keyframes $\{K_{i},K_{i+1}\}$ and their intermediate frames, we design a bidirectional tracking scheme to extract dense motion trajectories with improved temporal consistency. For each clip $\{F_{0},F_{1},...,F_{L-1}\}$ with temporal length $L$, we initialize isometric grids (grid size $N_{grid}\times N_{grid}$) in the first frame and employ Co-Tracker (karaev2024cotracker) to trace pixel-wise displacements originating from the grid vertices. The generated forward trajectory sequences over $L$ frames are represented as temporal chains of 3D coordinate points:
@@
To mitigate trajectory loss caused by occlusions or rotation, we reverse the temporal order of each clip and extract the corresponding backward trajectory sequences $\mathcal{T}_{bwd}$ using the same procedure. Subsequently, the backward trajectories are temporally realigned and concatenated with their forward counterparts to synthesize the final dense trajectories, which holistically preserve the motion characteristics between keyframes:
@@
To differentiate between distinct motion patterns, the raw trajectories are clustered using HDBSCAN (mcinnes2017hdbscan), a hierarchical density-based algorithm particularly effective in handling variable cluster densities and suppressing noise, which is critical for processing real-world videos with irregular motions. In this stage, $T_{i}$ is first transformed into 2D trajectories by excluding the points where $V_{t}^{i\ }=0$. Then, we formulate each $T_{i}$ as a spatio-temporal feature vector according to their coordinates. By clustering these features, the corresponding trajectories are adaptively categorized into different motion instances $\mathcal{T}^{0},\mathcal{T}^{1},...,\mathcal{T}^{n-1}$ and form the corresponding motion masks, as illustrated in Figure 2.
@@
The extracted dense motion trajectories inherently encapsulate the spatio-temporal dynamics of the source video, enabling the decoder to reconstruct inter-frame semantic relationships through diffusion-based generation (Section 3.3). However, noisy regions may be mistakenly identified as motion instances due to inaccuracies in motion clustering. Additionally, naively encoding these raw trajectories introduces significant bitrate overhead. To mitigate the influence of the accuracy of motion clustering and reduce inter-frame bitrate while preserving semantically critical motion information, we propose a trajectory sampling strategy driven by semantic loss awareness.
@@
We select the motion whose inter-frame semantic importance exceeds an empirically chosen threshold and extract sparse trajectories $\tau^{i}$ through k-means clustering, where the keypoint quantity $K$ for each instance is determined by the semantic importance score $S^{i}_{inter}$ and the quantity of trajectories $N^{i}$: